Find the values of x in the following equation using the rule that if a × b = 0 then a = 0
or b = 0. Do not expand the equation or use the quadratic formula.
(2x − 5)(3x + 1/5) = 0

Answer by mananth(16946)


(2x − 5)(3x + 1/5) = 0
therefore,
either (2x-5)=0 or (3x+1/5) =0
(2x-5)=0
2x=5
x=5/2

(3x + 1/5) = 0
3x=-1/5
x=-1/15